. A steel tank contains CO2 at 300K and a pressure of 10 atm calculate the internal
gas pressure when the tank gas
1
Expert's answer
2020-12-22T04:32:47-0500

We need to calculate the internal gas pressure when the gas tank is heated or cooled (not provided in the question). Let's say that final temperature of the tank and its contents is Tfinal.

p*V = n*R*T, R is the gas constant and V is constant because gas tank volume is constant

then p/T is constant

then pinitial/Tinitial = pfinal/Tfinal

then 10/300 = pfinal/Tfinal

and pfinal = Tfinal*10/300, if we know Tfinal we will know pfinal.